Write the slope-intercept equation for the line that passes through (10, 12) and is perpendicular to 5x + 8y = 1 Please show all of your work.

Answer by nerdybill(7384)

Write the slope-intercept equation for the line that passes through (10, 12) and is perpendicular to 5x + 8y = 1
.
Determine slope of:
5x + 8y = 1
8y = -5x + 1
y = (-5/8)x + 1/8
Slope is -5/8
.
New line, since it is perpendicular must have a slope that is the "negative reciprocal": 8/5
.
plug slope and given point into "point-slope" form:
y - y1 = m(x - x1)
y - 12 = (8/5)(x - 10)
y - 12 = (8/5)x - 16
y = (8/5)x - 4 (this is what they're looking for)